OpenStudy (anonymous):

I know i might be asking a bunch of questions but im just trying to refresh my mind for my finals tomorrow morning x-8/2x + 4 = x+3/x

OpenStudy (aravindg):

\[\huge \text{WELCOME TO OPENSTUDY!}\]

OpenStudy (anonymous):

Thanks

OpenStudy (aravindg):

oops .sorry .werll first add the like terms

OpenStudy (anonymous):

\[\frac{x-8}{2x + 4} = \frac{x+3}{x}\]like that?

OpenStudy (anonymous):

it is not really clear from your question, without any parentheses

OpenStudy (anonymous):

No the 4 is by itself

OpenStudy (aravindg):